OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. Write your solutions as a set of ordered pairs. 2x + 3y = 7

OpenStudy (anonymous):

x = {-1, 0, 2} 2x+3y=7 2(-1)+3y=7 -2+3y=7 3y=9 y=3 (-1,3) 2x+3y=7 2(0)+3y=7 3y=7 y=7/3 (0,7/3) 2x+3y=7 2(2)+3y=7 4+3y=7 3y=3 y=1 (2,1)